# An Adaptive Algorithm for Bilevel Optimization on Riemannian Manifolds

### v2

@@ Abstract
Existing methods for solving Riemannian bilevel optimization (RBO) problems require prior knowledge of the problem’s first- and second-order information and curvature parameter of the Riemannian manifold to determine step sizes, which poses practical limitations when these parameters are unknown or computationally infeasible to obtain. In this paper, we introduce the Adaptive Riemannian Hypergradient Descent (AdaRHD) algorithm for solving RBO problems. To our knowledge, AdaRHD is the first method to incorporate a fully adaptive step size strategy that eliminates the need for problem-specific parameters in RBO problem resolution. We prove that AdaRHD achieves an $\mathcal{O}(1/\epsilon)$ iteration complexity for finding an $\epsilon$-stationary point, thus matching the complexity of existing non-adaptive methods. Furthermore, we demonstrate that substituting exponential mappings with retraction mappings maintains the same complexity bound. Experiments demonstrate that AdaRHD achieves comparable performance to existing non-adaptive approaches while exhibiting greater robustness.
Key words: Adaptive algorithm, Bilevel optimization, Riemannian manifolds, Complexity
@@ 1 Introduction
Bilevel optimization has garnered significant attention owing to its diverse applications in fields such as reinforcement learning [44, 31], meta-learning [5, 60, 37], hyperparameter optimization [22, 63, 83], adversarial learning [6, 75, 76], and signal processing [45, 20]. This paper focuses on “Riemannian bilevel optimization (RBO)” [29, 19, 48], a framework that arises in applications such as Riemannian meta-learning [68], neural architecture search [65], image segmentation [24], Riemannian min-max optimization [40, 35, 88, 27, 77, 34], and Low-rank adaption [85]. The general formulation of RBO is as follows:
@@
where ${\mathcal{M}}_{x},{\mathcal{M}}_{y}$ are $d^{x}$- and $d^{y}$-dimensional complete Riemannian manifolds [48], the functions $f,g:{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}\rightarrow{\mathbb{R}}$ are smooth, the lower-level function $g(x,y)$ is geodesic strongly convex w.r.t. $y$, and the upper-level function $f$ is not required to be convex.
To address Problem (1), we extend the recent advance in Euclidean adaptive bilevel optimization [81] to propose adaptive Riemannian hypergradient methods. Although certain aspects of Euclidean analysis extend naturally to Riemannian settings, the manifold’s curvature parameters introduce distortions that pose unique analytical challenges. We identify key technical difficulties in designing adaptive methods for solving RBO problems as follows:
@@
    RBO necessitates that variables are constrained on Riemannian manifolds, making it intrinsically more complex than its Euclidean counterpart. Furthermore, critical parameters such as geodesic strong convexity, Lipschitz continuity, and curvature parameters are computationally demanding to estimate over manifolds. Consequently, the geometric structure of the manifold introduces distortions that create distinct analytical challenges in the convergence analysis of adaptive algorithms.
@@
    Compared to single-level Riemannian optimization, RBO inherently involves interdependent variable updates, resulting in complex step size selection challenges arising from coupled (hyper)gradient dynamics.
@@
    Unlike non-adaptive RBO methods, adaptive step size strategies may exhibit divergent behavior during initial iterations, which must be rigorously addressed to enable robust theoretical convergence analysis.
In this work, we present a comprehensive convergence analysis, demonstrating our approach to resolving these challenges.
@@ 1.1 Related works
In Table 1, we summarize relevant studies related to our method, providing their applicable scenarios, adaptive or not, and the computational complexity of first- and second-order information. Note that all constants (e.g., condition numbers) are omitted from the computational complexity analysis for simplicity. For details of the related works, please refer to Appendix A.
Table 1: Comparisons of first-order and second-order complexities for reaching $\epsilon$-stationary point. Here, “Euc” and “Rie” represent the feasible region of the algorithms are Euclidean space and Riemannian manifold, respectively. The notations “Det”, “F-S”, and “Sto” represent the applicable problems as deterministic, function-sum, and stochastic, respectively. Additionally, $G_{f}$ and $G_{g}$ are the gradient complexities of $f$ and $g$, respectively. $JV_{g}$ and $HV_{g}$ are the complexities of computing the second-order cross derivative and Hessian-vector product of $g$. The notation $\tilde{\mathcal{O}}$ denotes the omission of logarithmic terms in contrast to the standard $\mathcal{O}$ notation. Furthermore, the notation “NA” indicates that the corresponding complexity is not applicable.
@@
D-TFBO [81] Euc $\bm{\mathcal{\checkmark}}$ Det $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$
S-TFBO [81] $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-HINV [29] Rie Det $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ NA
RHGD-CG [29] $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-NS [29] $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-AD [29] $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RSHGD-HINV [29] F-S $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ NA
RieBO [48] Rie Det $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RieSBO [48] Sto $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$
RF2SA [19] Rie Det $\tilde{\mathcal{O}}(1/\epsilon^{3/2})$ $\tilde{\mathcal{O}}(1/\epsilon^{3/2})$ NA NA
RF2SA [19] Sto $\tilde{\mathcal{O}}(1/\epsilon^{7/2})$ $\tilde{\mathcal{O}}(1/\epsilon^{7/2})$ NA NA
AdaRHD-GD (Ours) Rie $\bm{\mathcal{\checkmark}}$ Det $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$
@@ 1.2 Contributions
    We introduce two strategies, including Riemannian gradient descent and tangent space conjugate gradient, to compute the approximate Riemannian hypergradient and establish error bounds between the exact and approximate hypergradients;
@@
    For solving smooth and strongly convex problems on Riemannian manifolds (the lower-level problem of Problem (1)), our proposed method is the first adaptive Riemannian gradient descent algorithm with a non-asymptotic convergence rate to our knowledge.
@@ 2 Preliminaries
In this section, we review standard definitions and preliminary results on Riemannian manifolds. All results presented here are available in the literature [47, 7, 29, 48], we restate them for conciseness.
We first establish essential properties and notation for Riemannian manifolds. The definition of a Riemannian manifold has been conducted in the literature [47, 7]. The Riemannian metric on a Riemannian manifold ${\mathcal{M}}$ at $x\in{\mathcal{M}}$ is denoted by $\langle\cdot,\cdot\rangle_{x}:T_{x}{\mathcal{M}}\times T_{x}{\mathcal{M}}\rightarrow{\mathbb{R}}$, with the induced norm on the tangent space $T_{x}{\mathcal{M}}$ defined as $\|u\|_{x}=\sqrt{\langle u,u\rangle_{x}}$ for an $u\in T_{x}{\mathcal{M}}$. We then recall the definitions of exponential mapping and parallel transport. For an tangent vector $u\in T_{x}{\mathcal{M}}$ and a geodesic $c:[0,1]\rightarrow{\mathcal{M}}$ satisfying $c(0)=x$, $c(1)=y$, and $c^{\prime}(0)=u$, the exponential mapping ${\rm Exp}_{x}:T_{x}{\mathcal{M}}\rightarrow{\mathcal{M}}$ is defined as ${\rm Exp}_{x}(u)=y$. The inverse of the exponential mapping, ${\rm Exp}_{x}^{-1}:{\mathcal{M}}\rightarrow T_{x}{\mathcal{M}}$, is called logarithm mapping [7, Definition 10.20], and the Riemannian distance between two points $x,y\in{\mathcal{M}}$ is given by $d(x,y)=\|{\rm Exp}_{x}^{-1}(y)\|_{x}=\|{\rm Exp}_{y}^{-1}(x)\|_{y}$. Parallel transport ${\mathcal{P}}_{x_{1}}^{x_{2}}:T_{x_{1}}{\mathcal{M}}\rightarrow T_{x_{2}}{\mathcal{M}}$ is a linear operator that preserves the inner product structure, satisfying $\langle u,v\rangle_{x_{1}}=\langle{\mathcal{P}}_{x_{1}}^{x_{2}}u,{\mathcal{P}}_{x_{1}}^{x_{2}}v\rangle_{x_{2}}$ for all $u,v\in T_{x_{1}}{\mathcal{M}}$.
@@
where ${\mathrm{D}}f(x)[u]$ denotes the directional derivative of $f$ at $x$ along $u$ [72, 7, 48]. For twice-differentiable $f$, the Riemannian Hessian ${\mathcal{H}}f(x)$ is defined as the covariant derivative of ${\mathcal{G}}f(x)$. A function $f\colon{\mathcal{M}}\to{\mathbb{R}}$ is “geodesically (strongly) convex” if $f(c(t))$ is (strongly) convex with respect to $t\in[0,1]$ for all geodesics $c\colon[0,1]\to\Omega$, where $\Omega\subseteq{\mathcal{M}}$ is a geodesically convex set [48, Definition 4]. Specifically, if $f$ is twice-differentiable, $\mu$-geodesic strong convexity of $f$ is equivalent to ${\mathcal{H}}f(x)\succeq\mu\mathrm{Id}$, where $\mathrm{Id}$ is the identity operator. Similarly, $f$ is “geodesically $L_{f}$-Lipschitz smooth” if
@@
Finally, for a bi-function $f\colon{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}\to{\mathbb{R}}$, the Riemannian gradients of $f$ w.r.t. $x$ and $y$ are denoted by ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$, respectively, while the Riemannian Hessians are ${\mathcal{H}}_{x}f(x,y)$ and ${\mathcal{H}}_{y}f(x,y)$. Moreover, the Riemannian cross-derivatives ${\mathcal{G}}_{xy}^{2}f(x,y)\colon T_{y}{\mathcal{M}}_{y}\to T_{x}{\mathcal{M}}_{x}$ and ${\mathcal{G}}_{yx}^{2}f(x,y)\colon T_{x}{\mathcal{M}}_{x}\to T_{y}{\mathcal{M}}_{y}$, discussed in [29, 48], are linear operators. Furthermore, the operator norm of a linear operator ${\mathcal{G}}\colon T_{y}{\mathcal{M}}_{y}\to T_{x}{\mathcal{M}}_{x}$ is defined as
@@
Building on the concepts of Lipschitz continuity (cf. Definition B.1) and geodesic strong convexity for functions defined on a Riemannian manifold, we present the following proposition.
Proposition 2.1 ([7, 29, 48]).
@@
Then, for the Lipschitzness of the functions and operators, and an important result of the trigonometric distance bound over the Riemannian manifolds, we present them in Appendix B.

### v3

@@ Abstract
Existing methods for solving Riemannian bilevel optimization (RBO) problems require prior knowledge of the problem’s first- and second-order information and curvature parameter of the Riemannian manifold to determine step sizes, which poses practical limitations when these parameters are unknown or computationally infeasible to obtain. In this paper, we introduce the Adaptive Riemannian Hypergradient Descent (AdaRHD) algorithm for solving RBO problems. To our knowledge, AdaRHD is the first method to incorporate a fully adaptive step size strategy that eliminates the need for problem-specific parameters in RBO. We prove that AdaRHD achieves an $\mathcal{O}(1/\epsilon)$ iteration complexity for finding an $\epsilon$-stationary point, thus matching the complexity of existing non-adaptive methods. Furthermore, we demonstrate that substituting exponential mappings with retraction mappings maintains the same complexity bound. Experiments demonstrate that AdaRHD achieves comparable performance to existing non-adaptive approaches while exhibiting greater robustness.
[…]
@@ 1 Introduction
Bilevel optimization has garnered significant attention owing to its diverse applications in fields such as reinforcement learning (konda1999actor, ; hong2023two, ), meta-learning (bertinetto2018meta, ; rajeswaran2019meta, ; ji2020convergence, ), hyperparameter optimization (franceschi2018bilevel, ; shaban2019truncated, ; yu2020hyper, ), adversarial learning (bishop2020optimal, ; wang2021fast, ; wang2022solving, ), and signal processing (kunapuli2008classification, ; flamary2014learning, ). This paper focuses on “Riemannian bilevel optimization (RBO)” (han2024framework, ; dutta2024riemannian, ; li2025riemannian, ), a framework that arises in applications such as Riemannian meta-learning (tabealhojeh2023rmaml, ), neural architecture search (sukthanker2020neural, ), image segmentation (ghadimi2018approximation, ), Riemannian min-max optimization (jordan2022first, ; huang2023gradient, ; zhang2023sion, ; han2023nonconvex, ; wang2023riemannian, ; hu2024extragradient, ), and low-rank adaption (zangrando2025debora, ). The general formulation of RBO is as follows:
@@
where ${\mathcal{M}}_{x},{\mathcal{M}}_{y}$ are $d^{x}$- and $d^{y}$-dimensional complete Riemannian manifolds (li2025riemannian, ), the functions $f,g:{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}\rightarrow{\mathbb{R}}$ are smooth, the lower-level function $g(x,y)$ is geodesic strongly convex w.r.t. $y$, and the upper-level function $f$ is not required to be convex.
To address Problem (1), we extend the recent advance in Euclidean adaptive bilevel optimization (yang2024tuning, ) to propose an adaptive Riemannian hypergradient method. Although certain aspects of Euclidean analysis extend naturally to Riemannian settings, the manifold’s curvature parameters introduce many analytical challenges. We propose key technical difficulties in designing adaptive methods for solving RBO problems as follows:
@@
    RBO necessitates that variables are constrained on Riemannian manifolds, making it intrinsically more complex than its Euclidean counterpart. Furthermore, critical parameters such as geodesic strong convexity, Lipschitz continuity, and curvature parameters over Riemannian manifolds are more computationally demanding to estimate than their Euclidean counterparts. Consequently, the geometric structure of the manifold creates distinct analytical challenges in the convergence analysis of adaptive algorithms.
@@
    Compared to single-level Riemannian (adaptive) optimization, RBO inherently involves interdependent variable updates, resulting in complex step size selection challenges arising from coupled (hyper)gradient dynamics.
@@
    Unlike non-adaptive RBO methods, adaptive step-size strategies can cause divergent behavior during initial iterations since the initial step size may be too large, which must be rigorously addressed to enable robust theoretical convergence analysis.
In this work, we present a comprehensive convergence analysis that demonstrates our approach to resolving these challenges.
@@ 1.1 Related works
Table 1 summarizes key studies on (Riemannian) bilevel optimization that are relevant to our methods. The table compares their applicable scenarios, whether they are adaptive, and their computational complexity of first- and second-order information. Note that constants, such as condition numbers, are omitted from the complexity analysis for simplicity. For a comprehensive review of the related works, please refer to Appendix A.
Table 1: Comparisons of first-order and second-order complexities for reaching an $\epsilon$-stationary point. Here, “Euc” and “Rie” represent the feasible region of the algorithms are Euclidean space and Riemannian manifold, respectively. The notations “Det”, “F-S”, and “Sto” represent the applicable problems as deterministic, function-sum, and stochastic, respectively. Additionally, $G_{f}$ and $G_{g}$ are the gradient complexities of $f$ and $g$, respectively. $JV_{g}$ and $HV_{g}$ are the complexities of computing the second-order cross derivative and Hessian-vector product of $g$. The notation $\tilde{\mathcal{O}}$ denotes the omission of logarithmic terms in contrast to the standard $\mathcal{O}$ notation. Furthermore, the notation “NA” represents that the corresponding complexity is not applicable.
@@
[…]
S-TFBO (yang2024tuning, ) $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-HINV (han2024framework, ) Rie Det $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ NA
RHGD-CG (han2024framework, ) $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-NS (han2024framework, ) $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-AD (han2024framework, ) $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RSHGD-HINV (han2024framework, ) F-S $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ NA
RieBO (li2025riemannian, ) Rie Det $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RieSBO (li2025riemannian, ) Sto $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$
RF2SA (dutta2024riemannian, ) Rie Det $\tilde{\mathcal{O}}(1/\epsilon^{3/2})$ $\tilde{\mathcal{O}}(1/\epsilon^{3/2})$ NA NA
RF2SA (dutta2024riemannian, ) Sto $\tilde{\mathcal{O}}(1/\epsilon^{7/2})$ $\tilde{\mathcal{O}}(1/\epsilon^{7/2})$ NA NA
[…]
@@ 1.2 Contributions
    We establish upper bounds for the total iterations of resolutions of the lower-level problem in Problem (1) and the corresponding linear system required for computing the Riemannian hypergradient. The derived bounds match the iteration complexity of the standard AdaGrad-Norm method for solving the strongly convex problems (xie2020linear, ).
@@
    We evaluate our algorithm against existing methods on various problems over Riemannian manifolds, including the Riemannian hyper-representation and robust optimization problems. The results demonstrate that our method performs comparably with non-adaptive methods while exhibiting significantly greater robustness.
@@ 2 Preliminaries
In this section, we review standard definitions and preliminary results on Riemannian optimization. All results presented here are available in the literature (lee2006riemannian, ; boumal2023introduction, ; han2024framework, ; li2025riemannian, ), we restate them for conciseness.
We first establish essential properties and notations for Riemannian manifolds. The definition of a Riemannian manifold has been conducted in the literature (lee2006riemannian, ; boumal2023introduction, ). The Riemannian metric on a Riemannian manifold $\mathcal{M}$} at $x\in{\mathcal{M}}$ is denoted by $\langle\cdot,\cdot\rangle_{x}:T_{x}{\mathcal{M}}\times T_{x}{\mathcal{M}}\rightarrow{\mathbb{R}}$, with the induced norm on the tangent space $T_{x}{\mathcal{M}}$ defined as $\|u\|_{x}=\sqrt{\langle u,u\rangle_{x}}$ for any $u\in T_{x}{\mathcal{M}}$. We then recall the definitions of exponential mapping and parallel transport. For an tangent vector $u\in T_{x}{\mathcal{M}}$ and a geodesic $c:[0,1]\rightarrow{\mathcal{M}}$ satisfying $c(0)=x$, $c(1)=y$, and $c^{\prime}(0)=u$, the exponential mapping ${\rm Exp}_{x}:T_{x}{\mathcal{M}}\rightarrow{\mathcal{M}}$ is defined as ${\rm Exp}_{x}(u)=y$. The inverse of the exponential mapping, ${\rm Exp}_{x}^{-1}:{\mathcal{M}}\rightarrow T_{x}{\mathcal{M}}$, is called logarithm mapping (boumal2023introduction, , Definition 10.20), and the Riemannian distance between two points $x,y\in{\mathcal{M}}$ is given by $d(x,y)=\|{\rm Exp}_{x}^{-1}(y)\|_{x}=\|{\rm Exp}_{y}^{-1}(x)\|_{y}$. Parallel transport ${\mathcal{P}}_{x_{1}}^{x_{2}}:T_{x_{1}}{\mathcal{M}}\rightarrow T_{x_{2}}{\mathcal{M}}$ is a linear operator that preserves the inner product structure, satisfying $\langle u,v\rangle_{x_{1}}=\langle{\mathcal{P}}_{x_{1}}^{x_{2}}u,{\mathcal{P}}_{x_{1}}^{x_{2}}v\rangle_{x_{2}}$ for all $u,v\in T_{x_{1}}{\mathcal{M}}$.
@@
where ${\mathrm{D}}f(x)[u]$ denotes the directional derivative of $f$ at $x$ along $u$ (tu2011manifolds, ; boumal2023introduction, ; li2025riemannian, ). For twice-differentiable $f$, the Riemannian Hessian ${\mathcal{H}}f(x)$ is defined as the covariant derivative of ${\mathcal{G}}f(x)$. A function $f\colon{\mathcal{M}}\to{\mathbb{R}}$ is “geodesically (strongly) convex” if $f(c(t))$ is (strongly) convex w.r.t. $t\in[0,1]$ for all geodesics $c\colon[0,1]\to\Omega$, where $\Omega\subseteq{\mathcal{M}}$ is a geodesically convex set (li2025riemannian, , Definition 4). Particularly, if $f$ is twice-differentiable, $\mu$-geodesic strong convexity of $f$ is equivalent to ${\mathcal{H}}f(x)\succeq\mu\mathrm{Id}$, where $\mathrm{Id}$ is the identity operator. Further, $f$ is “geodesically $L_{f}$-Lipschitz smooth” if
@@
Finally, for a bi-function $f\colon{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}\to{\mathbb{R}}$, the Riemannian gradients of $f$ w.r.t. $x$ and $y$ are denoted by ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$, respectively, while the Riemannian Hessians are ${\mathcal{H}}_{x}f(x,y)$ and ${\mathcal{H}}_{y}f(x,y)$. Moreover, the Riemannian cross-derivatives ${\mathcal{G}}_{xy}^{2}f(x,y)\colon T_{y}{\mathcal{M}}_{y}\to T_{x}{\mathcal{M}}_{x}$ and ${\mathcal{G}}_{yx}^{2}f(x,y)\colon T_{x}{\mathcal{M}}_{x}\to T_{y}{\mathcal{M}}_{y}$, discussed in han2024framework ; li2025riemannian , are linear operators. Furthermore, the operator norm of a linear operator ${\mathcal{G}}\colon T_{y}{\mathcal{M}}_{y}\to T_{x}{\mathcal{M}}_{x}$ is defined as
@@
Building on the concepts of Lipschitz continuity (cf. Definition B.1) and geodesic strong convexity for functions defined on Riemannian manifolds, we present the following proposition.
Proposition 2.1 (boumal2023introduction ; han2024framework ; li2025riemannian ).
@@
Additionally, for the Lipschitzness of the functions and operators, and an important result of the trigonometric distance bound over the Riemannian manifolds, we present them in Appendix B.